# Ceci N’est Pas un Drone: Investigating the Impact of Design Representation on Design Decision Making When Using GenAI

### v1

@@ 4.2 Study 2 - Engineering Students
Study 1 demonstrated that the number of options affects participants’ choices, and we observed a different trend in the problem with 8 design options compared to the problems with 2 or 16 options. To determine whether the irregularity is significantly influenced by the order in which the design problems were presented, we conducted the second study. The participants in Study 2A and Study 2B are engineering students at the aforementioned U.S. research university. The design problems for Study 2A are presented in the order of 2-option, 8-option, and 16-option, and those in Study 2B are presented in the order of 2-option, 16-option, and 8-option.
The results from Study 2A and Study 2B are first compared based on the accuracy of selecting the optimal designs. We performed two-way ANOVA tests on the design modality and the order in which the design questions are presented to see if the order of design questions would affect participants’ choices. The results of the two-way ANOVA tests are shown in Table 3 below. The order in which the design questions were presented did not significantly affect the accuracy of participants.
Design
Problem
Effect of Question Order Interaction Effect
2 options F(1, 375) = 0.172, p = 0.678 F(2, 375) = 0.322, p = 0.725
8 options F(1, 375) = 0.011, p = 0.917 F(2, 375) = 0.035, p = 0.965
16 options F(1, 375) = 2.235, p = 0.136 F(2, 375) = 1.924, p = 0.147
Table 3: Study 2: Two-way ANOVA test results on the interaction effect between the design modalities and the order in which the design questions are presented, for Study 2A and Study 2B. Results are based on the accuracy of participants in selecting the optimal designs.
Further, we performed an Extra‑Sum‑of‑Squares F–test on the participants’ accuracy to determine if the order of design questions affects participants’ overall responses. We compared two nested regression models, with the reduced model using "modality" and "question list" as predictors, and the full model using an additional predictor, "order of question list." The results suggest that there is no evidence that the additional variable (i.e., "order of question list") adds predictive power ($F(1,1135)=0.351,p=0.553$). The patterns of responses collected from Study 2A and Study 2B are not statistically different. Therefore, the results of Study 2A and Study 2B will be combined as one study and presented together. The number of participants that have chosen each design is shown in Figures 9, 10, and 11, with detailed distribution available in Appendix C, Tables 16, 17, and 18.
Figure 9: Study 2: Participants’ choice transition in Problem 1 - 2 options.
Figure 10: Study 2: Participants’ choice transition in Problem 2 - 8 options.
Figure 11: Study 2: Participants’ choice transition in Problem 3 - 16 options.
@@ 4.2.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem and for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 12 below. Overall, the entropy values are relatively small, indicating low randomness in the participants’ responses, suggesting the participants are making informed decisions based on the provided design information.
Figure 12: Study 2: Entropy of participants’ design choices
@@ 4.2.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 13. Most participants’ choices changed when the design modality changed from visual rendering to numerical performance data. Thus, design modalities did affect participants’ decision-making here. However, changes are inconsistent when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The changes in participants’ design choices are shown in Figures 9, 10, and 11.
@@
Participants’ selection of conventional designs (axis-symmetric and one-plane) versus unusual designs (non-axis-symmetric, non-planar, or off-plane) is shown in Figure 14. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked unusual designs when visual renderings became available (i.e., when transitioning from numerical performance data to the mixed modality).
@@ 4.2.3 Multiple Choice Question Response Accuracy
The average results of engineering student participants’ accuracy on selecting the optimal designs are shown in Figure 15, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is available in Appendix C, Table 19.
Figure 15: Study 2: Participants’ accuracy on selecting the optimal designs. Error bars represent 95% confidence interval. Participants are STEM students.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem 1 with 2 options: $F(2,378)=218.552$, $p<0.001$, $\eta_{p}^{2}=0.536$. Design problem 2 with 8 options: $F(2,378)=44.940$, $p<0.001$, $\eta_{p}^{2}=0.192$. Design problem 3 with 16 options: $F(2,378)=38.645$, $p<0.001$, $\eta_{p}^{2}=0.170$). The results indicate that the design modality would affect participants’ abilities to choose the "better" designs.
@@
Since different design problems result in different participants’ choices, we looked into the differences caused by the design modality separately for each design problem. Specifically, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 4. In general, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between numerical data and mixed modality conditions, and between visual rendering and mixed modality conditions.
@@
4.3 Study 3 - Aerospace Professionals
The participants in Study 3 are engineers and design professionals working in a large aerospace company. No statistical tests were performed on the data collected from Study 3, as the number of participants is notably smaller than what is required to satisfy the normality assumption of statistical tests. The number of participants that have chosen each design is shown in Figures 16, 17, and 18, with detailed distribution available in Appendix C, Tables  20, 21, and 22.
Figure 16: Study 3: Participants’ choice transition in Problem 1 - 2 options.
Figure 17: Study 3: Participants’ choice transition in Problem 2 - 8 options.
Figure 18: Study 3: Participants’ choice transition in Problem 3 - 16 options.
4.3.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 19 below. Overall, the entropy values are small, indicating low randomness in the participants’ answers, suggesting the participants are making informed decisions based on the provided design information.
Figure 19: Study 3: Entropy of participants’ design choices
4.3.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 20. Most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Interestingly, most participants’ design choices also changed when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. Design modalities did affect participants’ decision-making here. The transitions of participants’ design choice changes are shown in Figures  16, 17, and 18.
Figure 20: Study 3: Participants’ choice of design for consecutive questions in three design problems. V is with visual rendering, D is with numerical performance data, and M is with both visual rendering and numerical performance data.
Participants’ design selection of conventional design (axis-symmetric and one-plane) vs unusual design (non-axisymmetric, non-planar, or off-plane) is shown in Figure 21. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, when transitioning from numerical performance data to the mixed modality (both visual rendering and numerical performance data).
Figure 21: Study 3: Participants’ choice for conventional design vs unusual design in percentage.
4.3.3 Multiple Choice Question Response Accuracy
The average results of aerospace professional participants’ accuracy on selecting the optimal designs are shown in Figure 22, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is shown in Appendix C, Table 23.
Figure 22: Study 3: Participants’ accuracy on selecting the optimal designs. Error bars represent 95% confidence interval. Participants are aerospace professionals.
4.4 Textual Response
The frequencies of the driving factors for participants’ decision-making that appear in textual responses are shown in Figure 23, 24, and 25 below, for the three different design representation modality conditions.
Figure 23: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the visual rendering condition. The data is shown in the format of "number of responses (percentage)".
Figure 24: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the numerical performance data condition. The data is shown in the format of "number of responses (percentage)".
Figure 25: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the visual rendering + numerical performance data condition. The data is shown in the format of "number of responses (percentage)".
Overall, the results here demonstrated that the participants considered different factors when presented with different design representation modalities, and they frequently considered these driving factors when evaluating the designs. From the results of the frequencies of the driving factors for participants’ decision-making that appear in textual responses, we gained insight into the designers’ heuristics for their design choices.
When only visual renderings are provided, designers can effectively utilize their design heuristics and infer the real-world performance of the drone designs, mainly based on symmetry, stability, reliability, and cost of the design. Such a judgment is logical as symmetry, stability, and reliability can hint at the flight capabilities of the drone when no numerical performance data is provided. Interestingly, many participants in this study also estimated and compared the cost of the designs based on the visual rendering, a factor we thought only feasible with numerical performance data.
In comparison, when only numerical performance data is provided, participants focused on hover time, lift, and cost of the designs. These are also the factors that are most important for the flight task described in the design problem. Surprisingly, some participants also managed to infer information about stability, reliability of the design from the numerical data as well.
We also found that more participants considered the performance metrics than the visual inferences when both visual renderings and numerical performance data are available. The commonly considered design factors and features are similar to the previous two conditions, which include symmetry, stability, reliability, hover time, lift, and cost. This might suggest the participants in this study prioritize performance metrics over visual inferences when both are provided.
@@ 5.1 Design modalities affect design decision making
We found that different design modalities appear to affect engineers’ decision-making, and that only using a visual design rendering as a design modality can negatively impact engineers’ ability to identify the optimal design solutions, when using AI-powered generative design tools.
The results of the entropy tests suggest that all three tested modalities can provide some extent of design information, and participants are making informed design decisions, resulting in non-random selections and non-uniform distributions. From the distribution of participants’ answers, there are differences between showing visual design rendering, numerical performance data, and visual design rendering with numerical performance data. The design modality affects participants’ design decision-making process and leads to different results.
@@
The difference between different design modalities is more noticeable when looking at the participants’ choice of design for consecutive questions. For all three participant groups, most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Despite the inconsistency across design problems, a significant portion of participants’ design choices also changed when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. Since the design options are the same across all three modalities, the changes in participants’ choice when facing different design modalities indicate that the design modalities did affect participants’ decision-making. The participants change their minds and possibly select based on different factors when presented with different design modalities, even though the options are the same.
The difference between design modalities is also noticeable when looking at the participants’ ability to select the optimal designs. The three design modalities showed significantly different levels of benefits for selecting the Pareto-TOPSIS optimal design.
@@
Overall, the numerical performance data of the designs gives the most helpful and accurate information for finding the optimal design options. The participants were able to find the optimal designs with much higher accuracy when shown only numerical performance data, compared to the other two modalities. This also shows people’s ability to read and interpret design information from data and spider plots, even though the spider plots face some criticism, including inconsistent areas and shapes caused by axis ordering, misleading area size, and deceptive importance of irrelevant options due to dimension normalization, from the data visualization research community feldman_filled_2013; heijungs_two_2022; duan_origami_2023; abeynayake_efficacy_2023.
@@
The results also suggest that visual renderings of the designs alone might provide the least helpful design information for determining the optimality of the given designs. Participants’ accuracy in selecting optimal designs is quite low. It might suggest that the participants gained little information from the visual rendering of the design alone, or that the visual renderings alone provide distorted design information and lead the participants to inaccurate conclusions (e.g., the propellers can appear smaller if the body of the UAV is larger). However, an alternative explanation is that the participants gained additional information from the visual renderings that are not captured by the performance data and our simulator (e.g., a design might not be easily serviceable, or it might not perform well in cross-winds). Therefore, the difference in accuracy in selecting the optimal designs for the visual rendering condition might suggest that engineers and designers can utilize human heuristics and their design experiences to gain insights that are not easily captured by current computational tools.
@@
Interestingly, participant accuracy decreased when both the visual rendering and the numerical performance data were provided, compared to the numerical performance data only condition. It seems that the addition of visual information changed the participants’ minds, and they then made selections based on other factors. Participants’ textual responses suggest that they are considering and utilizing factors not explicitly shown with the numerical data. For example, in design problem 2 with 8 options, one of the aerospace professional participants (Study 3, P#01) in this study recognized the Pareto-TOPSIS optimal design based on the numerical performance data, but also pointed out the potential design flaws and infeasibility for real-life operations of that design. They then chose the designs with high reliability, stability, and practicality, which were factors they inferred from the visual rendering. Similar to the findings in the visual rendering condition, it is possible that the participants are getting additional information from the rendering that isn’t captured by the performance data. It could also mean the visual renderings are biasing people’s perceptions or causing design fixations. Future work should investigate further whether the addition of visual rendering is positively or negatively impacting designers’ decision-making.
@@ 5.2 Designers prefer conventional designs with good performance
In this study, participants showed clear preferences for conventional designs. Drone designs with conventional layouts and features have a significantly higher pick rate, namely, designs that are axis-symmetric and in which all propellers and the main body sit on one plane. Across all three participant pools, the participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, after transitioning from numerical performance data to the mixed modality.
@@
Moreover, there is a significant difference in pick rates for designs with conventional features (axisymmetric and one-plane) versus designs with atypical features (non-axisymmetric, non-planar, or off-plane) for the visual rendering only condition ($t(24)=3.103,p=0.014,\text{Cohen's }d=1.458$). Participants picked more conventional axisymmetric and one-plane designs significantly more often than the other designs. Furthermore, the axisymmetricity seems to be the dominant factor, as there is a significant difference in pick rates for axisymmetric designs versus non-axisymmetric designs for the visual rendering only condition ($t(24)=2.512,p=0.023,\text{Cohen's }d=0.866$). Participants prefer axisymmetric designs. In addition, we found that a Rod Length to Propeller Area Ratio of 5 is a critical divider for participants’ preferences. The pick rates for designs with Rod Length to Propeller Area Ratio greater than 5 are significantly higher than those with Rod Length to Propeller Area Ratio less than 5 ($t(24)=2.840,p=0.017,\text{Cohen's }d=1.209$).
@@
Looking at the changes in participants’ choice of design when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data, a noticeable number of participants moved away from the "unorthodox" designs with better performance, to pick designs with more conventional design features. An example is in design problem 1, where 55 out of 155 participants who chose the "unorthodox" design (design 18393) when shown only numerical performance data moved away and chose the more conventional design 16875 when presented with both visual rendering and numerical performance data (Figure 26). Similarly, 21 and 14 out of 38 and 66 participants, who chose the more unusual design 16763 and design 20985, two of the most picked designs, when provided only numerical performance data, chose the more conventional design 15317 instead when presented with both visual rendering and numerical performance data (Figure 27). These changes in participants’ choice of design when the design modality changed from numerical performance data to visual rendering + numerical performance data clearly show participants’ preference for conventional design features, despite the "unorthodox" designs having better performance data.
@@
Figure 26: Changes in participants’ choice of design for Design Problem 1 when the design modality changed from numerical performance data to visual rendering + numerical performance data.
@@
Figure 27: Changes in participants’ choice of design for Design Problem 3 when the design modality changed from numerical performance data to visual rendering + numerical performance data.
Among the designs with conventional features, design options with better performance have a higher pick rate when numerical performance data is provided. The participants’ accuracy in selecting the Pareto optimal designs within conventional designs is shown in Figures 28, 29, and 30 below (design problem 1 with two options is ruled out because it has only one conventional design option). Overall, the accuracy is high when numerical performance data is provided, in both numerical data and mixed modality conditions, suggesting that participants can identify the better-performing designs among those with conventional features.
Figure 28: Study 1: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are drone hobbyists.
Figure 29: Study 2: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are STEM students.
Figure 30: Study 3: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are aerospace professionals.
@@
Participants’ textual responses for their reasons for the design choices also reflect their preferences for axis-symmetric and "reasonable-looking" designs. One participant in Study 2A said they “feel that symmetrical designs have a big impact on the hover time, so asymmetrical design D, design E, and design H were immediately ruled out” [Study 2A, P#57]. A participant in Study 2A also ruled out asymmetrical designs as “options have large asymmetry would make maintenance harder and more expensive due to specialized parts”[Study 2A, P#19]. Participant #54 in Study 2B explained their design choice as “the rendering seems reasonable”. Another participant explained their preference for symmetrical design, as “symmetrical design would make manufacturing easier” [Study 2A, P#26]. Reliability can be another factor that drives the preference for symmetry, as one participant said “Design D combines symmetry with rotors close to the center of mass to create a reliable design” [Study 2A, P#18].
@@
Overall, participants prioritize designs with conventional features, and they prefer the better-performing designs among those with conventional features. Such design heuristics can be helpful for participants to identify the good performers, considering both simulation results and real-life scenarios. However, it can also mean that the participants are unknowingly or unintentionally omitting good-performing designs with atypical design features, leading to potential design fixation.
However, it is important to keep in mind that the performance data and the simulator used in this work may be inaccurate. Some real-life factors can be neglected, leading to an inaccurate evaluation of the drone design performance in real-life scenarios. The participants, however, may have considered these factors, which may have informed their design choice. This is further discussed in the limitations and future works subsection. Also, it is possible that the participants are aware of this imperfection during the studies, and therefore have less faith in the numerical performance data, leading to the preferences for conventional designs seen here.

### v2

@@ 4.2 Study 2 - Engineering Students
The participants in Study 2A and Study 2B are engineering students at the aforementioned U.S. research university. The design problems for Study 2A are presented in the order of 2-option, 8-option, and 16-option, and those in Study 2B are presented in the order of 2-option, 16-option, and 8-option.
The results from Study 2A and Study 2B are first compared based on the accuracy of selecting the optimal designs. We performed two-way ANOVA tests on the design modality and the order in which the design questions are presented to see if the order of design questions would affect participants’ choices (Design problem with 2 options: Effect of Question Order $F(1,375)=0.172,p=0.678$, Interaction Effect $F(2,375)=0.322,p=0.725$. Design problem with 8 options: Effect of Question Order $F(1,375)=0.011,p=0.917$, Interaction Effect $F(2,375)=0.035,p=0.965$. Design problem with 16 options: Effect of Question Order $F(1,375)=2.235,p=0.136$, Interaction Effect $F(2,375)=1.924,p=0.147$). The order in which the design questions were presented did not significantly affect the accuracy of participants.
Further, we performed an Extra‑Sum‑of‑Squares F–test on the participants’ accuracy to determine if the order of design questions affects participants’ overall responses. We compared two nested regression models, with the reduced model using "modality" and "question list" as predictors, and the full model using an additional predictor, "order of question list." The results suggest that there is no evidence that the additional variable (i.e., "order of question list") adds predictive power ($F(1,1135)=0.351,p=0.553$). The patterns of responses collected from Study 2A and Study 2B are not statistically different. Therefore, the results of Study 2A and Study 2B will be combined as one study and presented together. The number of participants that have chosen each design and the participants’ choice transitions are shown in Figures 13, 14, and 15 in Appendix C.
@@ 4.2.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem are calculated (Design problem with 2 options: $H_{max}=1$, $H_{Visual}=0.330$, $H_{Data}=0.112$, $H_{Visual+Data}=0.640$. Design problem with 8 options: $H_{max}=3$, $H_{Visual}=1.584$, $H_{Data}=1.445$, $H_{Visual+Data}=1.688$. Design problem with 16 options: $H_{max}=4$, $H_{Visual}=1.764$, $H_{Data}=1.744$, $H_{Visual+Data}=1.973$). Overall, the entropy values are relatively small, indicating low randomness in the participants’ responses, suggesting the participants are making informed decisions based on the provided design information.
@@ 4.2.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 5. Most participants’ choices changed when the design modality changed from visual rendering to numerical performance data. Thus, design modalities did affect participants’ decision-making here. However, changes are inconsistent when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The changes in participants’ design choices are shown in Figures 13, 14, and 15.
@@
Participants’ selection of conventional designs (axisymmetric and one-plane) versus unusual designs (non-axisymmetric, non-planar, or off-plane) is shown in Figure 6. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked unusual designs when visual renderings became available (i.e., when transitioning from numerical performance data to the mixed modality).
@@ 4.2.3 Multiple Choice Question Response Accuracy
The average results of engineering student participants’ accuracy on selecting the optimal designs are shown in Figure 7, with error bars indicating a 95% confidence interval.
Figure 7: Study 2: Participants’ accuracy on selecting the optimal designs. Error bars represent 95% confidence interval. p-value annotation legend: ns: 1.70e-02 < p <= 1.00e+00;    *: 1.00e-02 < p <= 1.70e-02;    **: 1.00e-03 < p <= 1.00e-02;    ***: 1.00e-04 < p <= 1.00e-03;    ****: p <= 1.00e-04. Participants are STEM students.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem with 2 options: $F(2,378)=218.552$, $p<0.001$, $\eta_{p}^{2}=0.536$. Design problem with 8 options: $F(2,378)=44.940$, $p<0.001$, $\eta_{p}^{2}=0.192$. Design problem with 16 options: $F(2,378)=38.645$, $p<0.001$, $\eta_{p}^{2}=0.170$). The results indicate that the design modality would affect participants’ abilities to choose the "better" designs.
@@
Further, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 3. In general, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between numerical data and mixed modality conditions, and between visual rendering and mixed modality conditions.
@@ 5.1 Design modalities affect design decision making
We found that different design modalities appear to affect engineers’ decision-making, and that using only visual design renderings as a design modality can limit engineers’ ability to identify the optimal design solutions when using AI-powered generative design tools.
@@
For both participant groups, most participants changed their design choices when the design modality switched from visual rendering to numerical performance data. Also, a significant portion of participants changed their design choices when the design modality switched from numerical performance data to the mixed modality. Since the design options remained the same, these changes indicate that design modalities did affect participants’ design choices. Participants changed their minds and possibly selected based on different factors when presented with different design modalities.
@@
The difference between design modalities is also evident in participants’ ability to select optimal designs. The three design modalities showed significantly different levels of benefits for selecting the optimal design. Participants identified optimal designs with much higher accuracy when shown only numerical performance data than when shown visual renderings or mixed modality. This result also demonstrates people’s ability to read and interpret design information from numerical data and spider plots, even though spider plots face criticism from the data visualization research community, including inconsistent areas and shapes caused by axis ordering, misleading area size, and deceptive importance of irrelevant options due to dimension normalization feldman_filled_2013; heijungs_two_2022; duan_origami_2023; abeynayake_efficacy_2023.
@@
Participants’ accuracy in selecting optimal designs was low when they saw only visual renderings. This might suggest that visual renderings alone either provide limited information about performance or convey distorted cues (e.g., propellers appearing smaller when the UAV body is larger). However, an alternative explanation is that participants gained additional information from the visual renderings that is not captured by the performance data or by our simulator (e.g., serviceability, vibration control, performance in cross-winds). In that case, the difference in accuracy in selecting optimal designs in the visual rendering condition might suggest that engineers and designers can leverage human heuristics and prior experiences to gain insights that are not easily captured by current computational tools.
@@
Interestingly, participants’ accuracy decreased when both the visual rendering and the numerical performance data were provided, compared to the numerical performance data only condition. It seems that the added visual information changed the participants’ minds, and they then made selections based on other factors. As in the visual rendering condition, participants may be getting additional information from the rendering that isn’t captured by the performance data. It could also mean the visual renderings are biasing people’s perceptions or causing design fixations.
@@ 5.2 Designers prefer conventional designs with good performance
In this study, participants showed clear preferences for conventional designs. Drone designs with conventional layouts and features have a significantly higher pick rate, namely, designs that are axisymmetric and in which all propellers and the main body lie on a single plane. Across both participant pools, participants showed a strong preference for conventional designs when provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, after transitioning from numerical performance data to the mixed modality.
@@
Moreover, there is a significant difference in pick rates for designs with conventional features (axisymmetric and one-plane) versus designs with atypical features (non-axisymmetric, non-planar, or off-plane) for the visual rendering only condition ($t(24)=3.103,p=0.014,\text{Cohen's }d=1.458$). Participants picked more conventional axisymmetric and one-plane designs significantly more often than the other designs. Furthermore, the axisymmetricity seems to be the dominant factor, as there is a significant difference in pick rates for axisymmetric designs versus non-axisymmetric designs for the visual rendering only condition ($t(24)=2.512,p=0.023,\text{Cohen's }d=0.866$). Participants prefer axisymmetric designs.
@@
Changes in design choices across modalities further illustrate this preference. When the design modality changed from numerical performance data to the mixed modality, a noticeable number of participants moved away from the "unorthodox" designs with better performance toward designs with more conventional design features. This phenomenon is best illustrated with the first design problem with 2 options, where 55 out of 155 participants who chose the "unorthodox" design (design 18393) when shown only numerical performance data moved away and switched to the more conventional design 16875 when presented with both visual rendering and numerical performance data (Figure 8). Similarly, in the third design problem with 16 options, 21 and 14 out of 38 and 66 participants, who chose the more unusual design 16763 and design 20985, two of the most picked designs, when provided only numerical performance data, switched to the more conventional design 15317 instead when presented with both visual rendering and numerical performance data (Figure 9). These transitions clearly show participants’ preference for conventional design features regardless of performance data.
@@
Figure 8: Changes in participants’ choice of design for Design Problem with 2 options when the design modality changed from numerical performance data to visual rendering + numerical performance data.
@@
Figure 9: Changes in participants’ choice of design for Design Problem with 16 options when the design modality changed from numerical performance data to visual rendering + numerical performance data.
@@
Participants’ textual responses for their reasons for the design choices also reflect their preferences for axisymmetric and "reasonable-looking" designs. One participant in Study 2A said they “feel that symmetrical designs have a big impact on the hover time, so asymmetrical design D, design E, and design H were immediately ruled out” [Study 2A, P#57]. A participant in Study 2A also ruled out asymmetrical designs as “options have large asymmetry would make maintenance harder and more expensive due to specialized parts”[Study 2A, P#19]. Participant #54 in Study 2B explained their design choice as “the rendering seems reasonable”. Another participant explained their preference for symmetrical design, as “symmetrical design would make manufacturing easier” [Study 2A, P#26]. Reliability can be another factor that drives the preference for symmetry, as one participant said “Design D combines symmetry with rotors close to the center of mass to create a reliable design” [Study 2A, P#18].
@@
Overall, participants prioritize designs with conventional features, and they prefer the better-performing designs among those with conventional features. Such design heuristics can be helpful for participants to identify the good performers, considering both simulation results and real-life scenarios. However, it can also mean that the participants are unknowingly or unintentionally omitting good-performing designs with atypical design features, leading to potential design fixation. Our work provides empirical evidence and insights into the prior work of Demirel et al. demirel_human-centered_2024, showing how a human-centered generative design framework can be improved by choosing appropriate representations to capture and incorporate human factors and preferences.